OpenStudy (tia_sweetheart17):

graph the linear equation. x-y=4

OpenStudy (anonymous):

so put it into slope intercept form by first adding y to both sides so you get: 4+y=x now subtracting 4 from both sides: y=x-4 so the slope intercept form is y=x-4

OpenStudy (anonymous):

do you know how to graph equations in slope intercept?

OpenStudy (anonymous):

so 1 is the slope and -4 in the y-intercept

OpenStudy (anonymous):

so you know the line crosses the y-axis at y=-4

OpenStudy (anonymous):

and its slope is 1 so for every 1 unit it goes to the right it goes 1 unit up

OpenStudy (anonymous):

so go 1 unit right of y=-4 and 1 unit up

OpenStudy (anonymous):

those r ur two point now connect them and make a line

OpenStudy (tia_sweetheart17):

thank u

OpenStudy (anonymous):

ok hope i helped

OpenStudy (tia_sweetheart17):

u did